Skip to Search

 
diagonal spacer

Systems Staff Meeting Notes
June 18, 2007

Present: Thomas Bielicke, Cynthia Bowker, Will Chaney, Logan Clapp, Debra Fetzer, Winston Harris, Bobbie Meng, Chris Nicolich, Amy Polk, Mark Sullivan, Stephen Williams
Announcements
  1. Bobbie has finished moving.
Review Minutes of 5/31/07
  1. Remove Winston Harris from the list of those in attendance.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Orders on hold until July 10.
  2. RAM received. It has not been installed yet. Possible procedures for installation discussed. Need to follow-up with DLC about the status of room 7B. It is a small area; the monitors would really crowd it. Also, need to get carriages to mount under the tables. We'll need about 70; they may have about half of that in Facilities. Need to check with Laurie and Michael in MSL about their machines, storage space and memory installation.
  3. ECCN - Export Control Classification Number - needs to be added to the note field in IRDB. Only applies to laptops at this time. ECCN should be obtained from Campus DSR.
  4. West security videos are now working.
  5. Xerox print services - PDF problem in West - Marco is working to correct it.
  6. SAN pieces - 3 UPS's received. NT2R to be replaced. Need to put together VM ware training.
Liaisons
  1. Vista: staff orientation announcements on hold.
Networking
  1. Web cluster failover - problem no longer occurring. Permissions set from cluster administration.
  2. Winston asks about regularly testing the Verify Patron failover. Amy asks that we wait until the new public image has been loaded.
  3. WSUS is set to not restart the servers. Some may need to be restarted manually.
  4. Server upgrades to W2K3 - all done except NT25.
  5. Upgrading Systems staff to VISTA - ready to go, but need to push it as an image. Logan will create a prototype so we can be sure the BDE is installed and runs correctly.
  6. Adding/removing machines from the domain - some problems are still occurring. Could be a problem with machines that Evan imaged - both public and staff.
  7. Bobbie spoke with Mary at CNS about the upgrade (Gigabit Ethernet) for East - it is on hold right now while they negotiate a campus-wide discount on parts. They are putting together a major package. It will take at least a month to receive the parts. They will upgrade over two weekends: the north tower one weekend and the south tower the next. Will probably happen sometime in August. If there is some critical need before that time, they may be able to do a temporary fix. IP phones (not soft phones) have been ordered for East.
  8. All of Journalism is upgrading this weekend, if the parts come in. Need to notify the Journalism library.
Software Projects Status
  1. Shadow system for payroll? On hold until further notice from Bill.
  2. Systems Management Tool has been named: GROVER (Grand Repository of Very Excellent Resources). We are not yet working on Web design; rather, we are creating routines to transfer information into the system so we will have real data to play with and to create foundation routines.
  3. Email set up for Michelle Foss and Scott Fagen to get a certificate approved from GeoTrust for ARES. It is to be sent to SYSHELP for approval at the time of renewal. When it comes in, assign it to Will.
  4. Jennifer Koontz - FCLA - sent a list of changes to Patron Update for v.18.
  5. Erich had some issues last week on things that had been approved and rolled out. Things will now be put on a test database for his group to evaluate prior to roll-out.
  6. Mark - Ephemeral Cities - presentation to grant agency on Sunday.
  7. Blog - still in testing.
Web Projects
  1. Will is the tentative back-up for Tom Minton.
Systems Security Management
  1. Meeting held.

 

University of Florida Home Page